OpenStudy (anonymous):

The value of 1/4+1/4.3+1/4.3^2+1/4.3^3=?

OpenStudy (anonymous):

1/4(1+1/3+1/3^2+1/3^3) 1/4(1-(1/3)^4)*1 ---------------- 1-1/3

OpenStudy (anonymous):

is this an infinite series or finite? that is \[\frac{1}{4}(1+\frac{1}{3}+\frac{1}{3^2} +\frac{1}{3^3}...)\]

OpenStudy (anonymous):

or is it just the one you wrote?

OpenStudy (anonymous):

could you plz explain how you get 1/3

OpenStudy (anonymous):

finite geometric series

OpenStudy (anonymous):

1/3 is the common ratio

OpenStudy (anonymous):

it is not \[\frac{1}{3}\] if it is just the terms you wrote.

OpenStudy (anonymous):

its 1/4 + 1/4.3, how it becomes 1+1/3

OpenStudy (anonymous):

1/4 is common

OpenStudy (anonymous):

oh is it 4.3 i took it 4*3???

OpenStudy (anonymous):

problem is odd if it is exactly what you wrote, then you need a calculator. can you post a picture of the original problem?

OpenStudy (anonymous):

Yes, its 4.3

OpenStudy (anonymous):

yeah, otherwise its simple one

OpenStudy (anonymous):

then satellite is right. u just need a calculator